OpenStudy (revircs):

How can I go about differentiating (3x^2)/((x-4)^(2/3)) ? I think I have the answer, but I am not completely sure.

OpenStudy (revircs):

For clarification: \[\frac{ 3x^2 }{ (x-4)^{2/3} }\]

OpenStudy (solomonzelman):

you can re-write it as: \(\large\color{black}{ \displaystyle (3x^2)((x-4)^{-2/3}) }\) and then apply the product rule.

OpenStudy (solomonzelman):

what is the derivative of: [1] \(\large\color{black}{ \displaystyle 3x^2 }\) [2] \(\large\color{black}{ \displaystyle (x-4)^{-2/3} }\)

OpenStudy (revircs):

1) 6x 2) \[-2/3(x-4)^{-5/3}\]

OpenStudy (revircs):

and after the product rule, it would be (6x) ((x-4)^(-2/3) + (3x^2)(-2/3(x-4)^(-5/3)

OpenStudy (solomonzelman):

yes, that is right!

OpenStudy (revircs):

simplified, it would be \[6x(x-4)^{-2/3}-\frac{ 6x^2 }{ 3(x-4)^{5/3} }\] is this the complete differentiation?